weigh less than 12 pounds? 150 and σ = 15 . Find the x in each of the following. 3. Suppose X is normal with μ a) P(X>x)-0.8765 b) PX<x 0.6330
Answer #1

Ans:

Given that

mean=150

standard deviation=15

a)P(Z>z)=0.8765

P(Z<=z)=1-0.8765=0.1235

z=-1.158

x=150-1.158*15=132.63

b)

P(Z<z)=0.6330

z=0.3398

x=150+0.3398*15=155.097